# Learning to Generate Parameters of ConvNets for Unseen Image Data

### v1

@@ IV-A Dataset Construction
In the experiment, we construct numerous datasets for evaluating our method based on several image datasets: Fashion-MNIST [14], CIFAR-100 [15], Mini-ImageNet [16], Animals-10 [56], CIFAR-10 [15], DTD [57], and a larger scale dataset, ImageNet-1K  [5]. The constructed datasets are based on two following settings: Intra-dataset prediction setting and Inter-dataset prediction setting.
The intra-dataset prediction setting is described as follows:
Fashion-set: We construct 2000 groups of datasets from the samples of 6 classes on Fashion-MNIST to train PudNet. The testing groups of datasets are constructed from samples of the remaining 4 classes. we construct 500 groups of datasets from the 4-category testing set to generate 500 groups of network parameters. For each group of network parameters, we also construct another dataset having identical labels but not having overlapped samples with the dataset used for generating parameters, in order to test the performance of the predicted network parameters.
CIFAR100-set: We choose 80 classes from CIFAR-100 for training sets and 20 not overlapped classes for testing sets. We sample 100000 groups of datasets for training. Similar to Fashion-set, we construct 500 groups of datasets for parameter prediction, and another 500 groups for testing.
ImageNet100-set: Similar to CIFAR100-set, we sample 50000 groups of datasets from 80-category training sets of Mini-ImageNet to train PudNet. We construct 500 groups of datasets from remaining 20-category testing sets to predict parameters, and sample another 500 groups for testing.
ImageNet1K-set: We partition ImageNet into mutually exclusive sets of 800 classes for training and 200 classes for testing. For the training set of 800 classes, we construct 20000 sub-datasets to train PudNet. We employ the testing data of 200 classes to evaluate the performance of PudNet.
Note that the training class set and testing class set in the above datasets are not overlapped.
To further verify our PudNet, we introduce an inter-dataset setting by constructing cross-domain datasets. We use ImageNet to construct training datasets. There are 50000 groups of datasets from ImageNet-set to train PudNet. The samples in Animals-10, CIFAR-10, and DTD are used for evaluation, respectively. Note that ImageNet and the above three other datasets are popular cross-domain datasets, which have been widely used for out-of-distribution (OOD) learning [58, 59], and cross-domain learning [60, 61]. The constructed datasets are summarized as follows:
@@
Method Fashion-set CIFAR100-set ImageNet100-set
@@ IV-B Baselines and Implementation Details
We compare our method with the most related work, i.e., traditional iterative based training paradigm including training from scratch with Adam [6] and one training acceleration method, GC [9]. We also take the pretrained model as a baseline. In addition, we also compare with meta-learning methods, including MatchNet [16], ProtoNet [40], Meta-Baseline [44], Meta-DeepDBC [45], and MUSML [43]. We use three kinds of architectures as our target network $\Omega$ respectively: a 3-layer CNN ConvNet-3, ResNet-18, ResNet-34. To ensure a fair comparison, we use all training datasets for training the related meta-learning methods and the pretrained model. For all experiments, we use the accuracy metric (ACC) metric to evaluate the classification performance. We mainly use the Top-1 Accuracy throughout the experiments, unless otherwise specified.
We perform the experiments using GeForce RTX 3090 Ti GPU. All experiments are optimized by the Adam optimizer. We set the learning rate to 0.001 and train PudNet until convergence. In metric-based learning process, following [44], the temperature $\tau=10$ in Eq.(7) is fixed. 10 labeled samples per class are used to deduce the class centroid. As mentioned before, we introduce auxiliary tasks to assist optimization. We add a full classification linear layer (e.g. 80-way linear head in CIFAR-100) to maintain static class set during training. We also introduce a consistency loss, since the dimension between logits deduced by metric-based classification (e.g.5-dimensional in CIFAR-100) and the logits produced by full linear head (e.g.80-dimensional in CIFAR-100) are not matched, we transpose the 5-dimensional logit to 80-dimensional logit, by padding the rest values with zero. We search $\eta$ from $\{0,0.1,0.2,0.3,\cdots,0.9\}$. For the target network ResNet-18, we set $\eta=0.2$ for Fashion-set, $\eta=0.5$ for CIFAR100-set and ImageNet-set. The hyperparameter $\eta$ is set as 0.5 on the inter-dataset setting datasets. For the target network ConvNet-3, we set $\eta=0.1$, and for the target network ResNet-34, we set $\eta=0.5$.
IV-C Result and Analysis
Performance Analysis of Intra-dataset Prediction. Table I shows the results of our method in the intra-dataset setting. We could find that our method consistently outperforms the meta-learning methods and the pretrained method. This demonstrates that learning a hyper-mapping between datasets and corresponding network parameters is effective. To demonstrate the time consumption our method could save, we also provide the time of training the model from scratch by a widely-used optimizer Adam and the training acceleration technique, GC. We find that it takes around 55, 119, 141 GPU seconds to train ResNet-18 using the accelerated method GC and the network obtains top-1 accuracies of 99.94%, 75.74%, 72.89% on Fashion-set, CIFAR-set, ImageNet-set respectively. While our method costs only around 0.5 GPU seconds to predict the parameters of ResNet-18 and still achieves a comparable performance (96.24%, 73.33%, 71.57% top-1 accuracies) on the three datasets respectively, at least 100 times faster than the accelerated method.
@@ IV-C Result and Analysis
TABLE IV: Results of different methods on ImageNet1K-set in the intra-dataset setting.
Method ImageNet-1K
@@
Meta-Baseline 38.92$\pm$0.93 -
Meta-DeepDBC 39.15$\pm$1.12 -
MUSML 41.33$\pm$0.95 35.17
@@
[…]
Scaling Up to ImageNet-1K. For the intra-dataset setting, we partition ImageNet-1K into mutually exclusive sets of 800 classes for training and 200 classes for testing. Top-5 accuracy of different methods are reported as in Table IV. We can observe that our method still achieves surprisingly good efficiency. For instance, It takes 8967 GPU seconds to train ResNet-18 on the ImageNet-1K using GC from scratch and obtain a top-5 accuracy of 44.65 %. However, our PudNet costs only 3.89 GPU seconds to predict the network parameters of ResNet-18 achieving comparable performance (44.92 %), more than 2,300 times faster than the traditional training paradigm. In addition, our method also outperforms the state-of-the-art meta-learning methods in a large margin. These show that PudNet can effectively scale to a larger and more complex dataset.
For the inter-dataset setting, we add another experiments on three inter-datasets: ImageNet-1K $\to$ Animals-10, ImageNet-1K $\to$ CIFAR-10 and ImageNet-1K $\to$ DTD. The results are listed as in Table V. Based on the above table, our method still achieves surprisingly good efficiency. Please kindly note that compared to our PudNet trained on ImageNet-100, as in Table III, which achieves 42.32 % accuracy in Animals-10, 45.07% accuracy in CIFAR-10, and 47.5% accuracy in DTD, training PudNet on the larger ImageNet dataset yields an improvement of approximately 16%, 6%, and 7% on the three inter-dataset tasks, respectively, while requiring comparable time costs. This indicates that training our model with more complex and comprehensive datasets results in better generalization performance. Based on the experiments on both intra-dataset and inter-dataset settings, our PudNet can effectively scale to larger datasets.
Ablation Study. We design several variants of our method to analyse the effect of different parts in both intra-dataset and inter-dataset settings. We first analyse the effect of the context relation information. The results are listed in Table VI. PudNet-w.o.-Context denotes our method directly feeds the dataset sketch into the weight generator without using AHRU. Our PudNet outperforms PudNet-w.o.-Context in a large margin, demonstrating the effectiveness of capturing dependencies among parameters of different layers.
TABLE VI: Ablation study of our method PudNet. ‘IN’ denotes ‘ImageNet’.
@@
TABLE VII: Results of different methods in terms of different target network architectures on CIFAR100-set.
@@
MatchNet 47.75$\pm$0.73 - 60.37$\pm$0.70 -
ProtoNet 51.96$\pm$0.57 - 64.28$\pm$0.58 -
Meta-baseline 57.69$\pm$0.38 - 67.40$\pm$0.69 -
Meta-DeepDBC 60.52$\pm$0.41 - 69.64$\pm$0.75 -
MUSML 56.49$\pm$0.56 1.21 66.39$\pm$0.59 3.11
Adam Scratch 1 epoch 49.11$\pm$1.03 0.99 47.39$\pm$1.36 5.47
@@
We further evaluate the effect of the auxiliary task. PudNet-metric means our method only uses the parameter-free loss. PudNet-w.o.-kl means our method does not use the KL Divergence. As shown in Table VI, PudNet-w.o.-kl has better performance than PudNet-metric, demonstrating the effectiveness of the auxiliary full classification task. PudNet outperforms PudNet-w.o.-kl, illustrating the effectiveness of matching the predicted probability distributions of the parameter-free method and full classification method.
Performance Analysis of Different Target Networks. We conduct experiments to further evaluate our method by predicting parameters of different target network architectures: ConvNet-3 and ResNet-34. The results are listed in Table VII. We observe that our method can obtain comparable accuracy to those of GC at 30 epochs for different target network architectures, while our method is more than 250 times faster than GC. This further demonstrates the efficiency of our method.
@@
MatchNet 56.83$\pm$0.40 52.26$\pm$0.71
ProtoNet 58.14$\pm$0.29 54.44$\pm$0.64
Meta-baseline 60.05$\pm$0.27 54.52$\pm$0.58
Meta-DeepDBC 61.35$\pm$0.28 55.09$\pm$0.63
MUSML 59.28$\pm$0.29 53.61$\pm$0.69
Adam Scratch 48.25$\pm$0.30 52.16$\pm$0.56
GC Scratch 50.03$\pm$0.32 53.58$\pm$0.61
@@
Effect of Different Groups of Datasets: We analyze the effect of different groups of datasets for training. Fig. 5 reports the results of exploiting PudNet to predict parameters for ResNet-18. We construct 2000, 100000, 100000 groups of datasets for training our PudNet on Fashion-set, CIFAR100-set, ImageNet-set respectively. We find that with more groups of datasets for training, our PudNet could obtain better performance. This is because with larger number of groups to learn the hyper-mapping relation, our PudNet could obtain better generalization ability. However, when the number of groups becomes large, the performance increase becomes slow.
@@
(b) CIFAR100-set with ResNet-18
(c) ImageNet-set with ResNet-18
@@
(b) metric loss on CIFAR100-set
(c) metric loss on ImageNet-set
(d) total train loss on Fashion-set
(e) total train loss on CIFAR100-set
(f) total train loss on ImageNet-set
@@
Parameter Sensitive Analysis: We analyze the effect of different values of the hyper-parameter $\eta$. Recall that $\eta$ controls the percent of dataset complementary information in the initial residual connection. Fig. 6(a)(b)(c) show the results in terms of ResNet-18. We observe that our model obtain better performance when $\eta>0$ in general. Additionally, our method is not sensitive to $\eta$ in a relatively large range.
@@
Convergence Analysis: We discuss the convergence property of the proposed method by plotting the loss curves with increasing iteration. Here we utilize PudNet to predict parameters for ResNet-18, based on Fashion-set, CIFAR100-set and ImageNet-set respectively. As shown in Fig. 7, the training metric-based loss (abbreviated as metric loss) and training total loss first decrease rapidly as the number of iterations increases, and then gradually decrease to convergence.
V Conclusion and Limitations

### v3

@@ IV-A Dataset Construction
In the experiment, we evaluate our method by constructing numerous datasets based on several image datasets: Fashion-MNIST [14], CIFAR-100 [15], Mini-ImageNet [16], Animals-10 [56], CIFAR-10 [15], DTD [57], and a large-scale dataset, ImageNet-1K  [5]. These constructed datasets are based on two distinct settings: the Intra-dataset prediction setting and the Inter-dataset prediction setting.
The intra-dataset prediction setting is described below:
Fashion-set: We construct 2000 groups of datasets from the samples of 6 classes in Fashion-MNIST to train PudNet. The testing groups of datasets are constructed from samples of the remaining 4 classes. We construct 500 groups of datasets from the 4-category testing set to generate 500 groups of network parameters. For each group of network parameters, we also construct another dataset having identical labels but not having overlapped samples with the dataset used for generating parameters, in order to test the performance of the predicted network parameters.
CIFAR-100-set: We select 80 classes from CIFAR-100 as the training sets and the remaining 20 non-overlapping classes as the testing sets. We randomly sample 100,000 groups of datasets for training. Similar to Fashion-set, we construct 500 groups of datasets for parameter prediction, and another 500 groups for testing.
ImageNet-100-set: Similar to CIFAR-100-set, we randomly sample 50,000 groups of datasets from 80-category training sets of Mini-ImageNet to train PudNet. We construct 500 groups of datasets from the remaining 20-category testing sets to predict parameters, and we sample another 500 groups of datasets for testing.
ImageNet-1K-set: We partition ImageNet-1K into two mutually exclusive sets of 800 classes for training and 200 classes for testing. Within the training set of 800 classes, we construct 20,000 sub-datasets to train PudNet. We employ the testing data of 200 classes to evaluate the performance of PudNet.
Please note that the training class set and testing class set in the aforementioned datasets are not overlapped.
To further validate the performance of our PudNet, we introduce an inter-dataset setting by constructing cross-domain datasets. We construct three training datasets: CIFAR-100-set with 100,000 dataset groups, ImageNet-100-set with 50,000 dataset groups, and ImageNet-1K-set with 20,000 dataset groups. Each of these datasets is used individually to train PudNet. Samples from Animals-10, CIFAR-10, and DTD are utilized for evaluation, respectively. Note that ImageNet and the above three datasets are popular cross-domain datasets, which have been widely utilized for out-of-distribution (OOD) learning [58, 59] and cross-domain learning [60, 61]. The summary of the constructed datasets is as follows:
@@
Method Fashion-set CIFAR-100-set ImageNet-100-set
@@ IV-B Baselines and Implementation Details
We compare our method with the most related work, i.e., traditional iterative based training paradigm such as training from scratch with Adam [6] and a training acceleration method, GC [9]. We also consider the pretrained model as a baseline for comparison. Furthermore, we compare our method with various meta-learning methods, namely MatchNet [16], ProtoNet [40], Meta-Baseline [44], Meta-DeepDBC [45], and MUSML [43]. For our target network $\Omega$, We explore three different architectures: a 3-layer CNN (ConvNet-3), ResNet-18[4], and ResNet-34[4]. Throughout the experiments, we utilize the accuracy (ACC) metric to evaluate the classification performance. Top-1 Accuracy is employed as the evaluation metric, unless otherwise specified.
We perform the experiments using GeForce RTX 3090 Ti GPU. All experiments are optimized by the Adam optimizer. We set the learning rate to 0.001 and train PudNet until convergence. In the metric-based learning process, following [44], the temperature $\tau=10$ in Eq.(7) is fixed. 10 labeled samples per class are used to deduce the class centroid. As mentioned before, we introduce auxiliary tasks to assist optimization. We incorporate a full classification linear layer (e.g. an 80-way linear head in CIFAR-100) to maintain a static class set during training. We also introduce a consistency loss to address the mismatch in dimensions between the logits derived from metric-based classification (e.g., 5-dimensional in CIFAR-100) and the logits produced by the full linear head (e.g., 80-dimensional in CIFAR-100). To align the dimensions, we transpose the 5-dimensional logits to the 80-dimensional logits by padding the remaining values with zeros. We search $\eta$ from $\{0,0.1,0.2,0.3,\cdots,0.9\}$. For the ResNet-18 target network, we set $\eta=0.2$ for Fashion-set, $\eta=0.5$ for CIFAR-100-set, ImageNet-100-set and ImageNet-1K-set. The hyperparameter $\eta$ is set as 0.5 in the inter-dataset setting. For the ConvNet-3 target network, we set $\eta=0.1$, and for the ResNet-34 target network, we set $\eta=0.5$.
@@
TABLE IV: Results of different methods on ImageNet-1K-set in the intra-dataset setting.
Method ImageNet-1K-set
@@
Meta-Baseline [44] 38.92$\pm$0.93 -
Meta-DeepDBC [45] 39.15$\pm$1.12 -
MH [63] 39.07$\pm$0.89 39.22
MUSML [43] 41.33$\pm$0.95 35.17
GAP [64] 42.57$\pm$0.87 36.21
@@
IV-C Result and Analysis
Performance Analysis of Intra-dataset Prediction. Table I shows the results of our method in the intra-dataset setting. We could find that our method consistently outperforms both the meta-learning methods and the pretrained method. This demonstrates the effectiveness of learning a hyper-mapping between datasets and their corresponding network parameters. To showcase the time savings achieved by our method, we provide the training time required for training the model from scratch using the widely-used optimizer, Adam, as well as the training acceleration technique, GC. We observe that it takes around 55, 119, and 141 GPU seconds to train ResNet-18 using the accelerated method GC and the network obtains top-1 accuracies of 99.94%, 75.74%, and 72.89% on Fashion-set, CIFAR-100-set, and ImageNet-100-set respectively. In contrast, our method only costs around 0.5 GPU seconds to predict the parameters of ResNet-18, while still achieving comparable performance (96.24%, 73.33%, and 71.57% top-1 accuracies) on the three datasets respectively, at least 100 times faster than the accelerated method.
@@ IV-C Result and Analysis
TABLE VI: Ablation study of our method PudNet. ‘IN’ denotes ‘ImageNet-100’.
@@
TABLE VII: Results of different methods in terms of different target network architectures on CIFAR-100-set.
@@
MatchNet [16] 47.75$\pm$0.73 - 60.37$\pm$0.70 -
ProtoNet [40] 51.96$\pm$0.57 - 64.28$\pm$0.58 -
Meta-baseline [44] 57.69$\pm$0.38 - 67.40$\pm$0.69 -
Meta-DeepDBC [45] 60.52$\pm$0.41 - 69.64$\pm$0.75 -
MUSML [43] 56.49$\pm$0.56 1.21 66.39$\pm$0.59 3.11
Adam Scratch [6] 1 epoch 49.11$\pm$1.03 0.99 47.39$\pm$1.36 5.47
@@
MatchNet [16] 56.83$\pm$0.40 52.26$\pm$0.71
ProtoNet [40] 58.14$\pm$0.29 54.44$\pm$0.64
Meta-baseline [44] 60.05$\pm$0.27 54.52$\pm$0.58
Meta-DeepDBC [45] 61.35$\pm$0.28 55.09$\pm$0.63
MUSML [43] 59.28$\pm$0.29 53.61$\pm$0.69
Adam Scratch [6] 48.25$\pm$0.30 52.16$\pm$0.56
GC Scratch [9] 50.03$\pm$0.32 53.58$\pm$0.61
@@
Scaling Up to ImageNet-1K. For the intra-dataset setting, we partition ImageNet-1K into two mutually exclusive sets of 800 classes for training and 200 classes for testing. The top-5 accuracies of different methods are reported in Table IV. We can observe that our method still achieves surprisingly good efficiency. For instance, It takes 8,967 GPU seconds to train ResNet-18 on ImageNet-1K using GC from scratch and obtains a top-5 accuracy of 44.65%. However, our PudNet only costs 3.89 GPU seconds to predict the network parameters of ResNet-18 achieving a comparable performance (44.92%), more than 2,300 times faster than the traditional training paradigm. In addition, our method also outperforms the state-of-the-art meta-learning methods by a large margin. These results highlight the effective scalability of PudNet to larger and more complex datasets.
For the inter-dataset setting, we conduct additional experiments on three cross-domain datasets: ImageNet-1K $\to$ Animals-10, ImageNet-1K $\to$ CIFAR-10, and ImageNet-1K $\to$ DTD. The results are presented in Table V. Remarkably, our method continues to demonstrate impressive efficiency. Please kindly note that compared to our PudNet trained on ImageNet-100, as in Table III, which achieves 42.43% top-1 accuracy on Animals-10, 45.07% top-1 accuracy on CIFAR-10, and 47.5% top-1 accuracy on DTD, training PudNet on the larger ImageNet dataset yields an improvement of approximately 16%, 6%, and 7% on the three inter-dataset tasks, respectively, while requiring comparable time costs. These results indicate that training our model with more complex and comprehensive datasets results in better generalization performance. Based on the experiments conducted in both the intra-dataset and inter-dataset settings, it is evident that our PudNet can effectively scale to larger datasets.
@@
(b) CIFAR-100-set with ResNet-18
(c) ImageNet-100-set with ResNet-18
@@
(b) metric loss on CIFAR-100-set
(c) metric loss on ImageNet-100-set
(d) total training loss on Fashion-set
(e) total training loss on CIFAR-100-set
(f) total training loss on ImageNet-100-set
@@
Ablation Study. We design several variants of our method to analyze the effect of different components in both intra-dataset and inter-dataset settings. The results are listed in Table VI. We first examine the effect of the context relation information. ‘PudNet-w.o.-Context’ denotes that the dataset sketch is directly fed into the weight generator without utilizing AHRU. Our PudNet outperforms ‘PudNet-w.o.-Context’ by a significant margin, demonstrating the effectiveness of capturing dependencies among parameters of different layers.
We further evaluate the effect of the auxiliary task. ‘PudNet-metric’ means our method only utilizes the parameter-free loss. ‘PudNet-w.o.-kl’ means our method does not use the KL Divergence. As shown in Table VI, ‘PudNet-w.o.-kl’ achieves better performance compared to ‘PudNet-metric’, highlighting the effectiveness of the auxiliary full classification task. Moreover, PudNet outperforms ‘PudNet-w.o.-kl’, thus illustrating the effectiveness of matching the predicted probability distributions of the parameter-free classification and full classification method.
Performance Analysis of Different Target Networks. We conduct experiments to further evaluate our method by predicting parameters for different target network architectures, namely ConvNet-3 and ResNet-34. The results are presented in Table VII. We observe that our method can achieve comparable top-1 accuracy to those of GC at 30 epochs for both target network architectures, while our method exhibits more than 250 times faster than GC. This further demonstrates the efficiency of our method.
TABLE IX: Results of different methods on ImageNet-1K-set in the Intra-dataset setting for image denoising
Method time(sec.) Gaussian ($\sigma=0.25$) Gaussian ($\sigma=0.50$) Gaussian ($\sigma=0.75$)
PNSR(dB) SSIM PNSR(dB) SSIM PNSR(dB) SSIM
Adam Scratch 1 epoch $0.13\times 10^{3}$ 21.69 0.871 20.88 0.851 18.98 0.814
50 epochs $6.72\times 10^{3}$ 28.01 0.947 25.84 0.919 24.21 0.891
100 epochs $1.36\times 10^{4}$ 28.45 0.951 26.25 0.920 24.65 0.893
300 epochs $4.17\times 10^{4}$ 29.56 0.956 26.78 0.927 25.02 0.902
GC Scratch 1 epoch $0.16\times 10^{3}$ 21.85 0.869 21.01 0.846 18.71 0.821
50 epochs $6.80\times 10^{3}$ 28.47 0.947 25.95 0.919 24.64 0.892
100 epochs $1.37\times 10^{4}$ 28.64 0.953 26.92 0.924 25.02 0.901
300 epochs $4.17\times 10^{4}$ 30.12 0.959 27.23 0.933 25.77 0.914
PudNet 0.91 28.53 0.950 25.97 0.921 24.52 0.892
TABLE X: Results of different methods in the Inter-dataset setting for image denoising
[…]
PNSR(dB) SSIM time(sec.) PNSR(dB) SSIM time(sec.) PNSR(dB) SSIM time(sec.)
Gaussian ($\sigma=0.25$) GC Scratch 1 epoch 21.52 0.891 $0.14\times 10^{3}$ 22.56 0.902 $0.11\times 10^{3}$ 21.04 0.723 $0.57\times 10^{2}$
50 epochs 27.28 0.957 $7.30\times 10^{3}$ 32.58 0.973 $5.34\times 10^{3}$ 26.18 0.809 $2.86\times 10^{3}$
100 epochs 28.87 0.963 $1.45\times 10^{4}$ 34.12 0.980 $1.08\times 10^{4}$ 27.43 0.834 $5.80\times 10^{3}$
300 epochs 29.51 0.968 $4.09\times 10^{4}$ 36.01 0.984 $3.20\times 10^{4}$ 29.01 0.851 $1.74\times 10^{4}$
PudNet 28.32 0.960 1.48 33.52 0.975 1.14 27.66 0.819 1.45
Gaussian ($\sigma=0.50$) GC Scratch 1 epoch 20.40 0.872 $0.14\times 10^{3}$ 21.65 0.887 $0.10\times 10^{3}$ 18.85 0.587 $0.57\times 10^{2}$
50 epochs 24.51 0.919 $7.30\times 10^{3}$ 30.53 0.961 $5.33\times 10^{3}$ 23.40 0.703 $2.87\times 10^{3}$
100 epochs 25.64 0.931 $1.45\times 10^{4}$ 31.28 0.964 $1.08\times 10^{4}$ 24.79 0.740 $5.81\times 10^{3}$
300 epochs 26.55 0.941 $4.09\times 10^{4}$ 32.14 0.971 $3.20\times 10^{4}$ 26.22 0.762 $1.74\times 10^{4}$
PudNet 25.49 0.930 1.49 30.82 0.960 1.16 25.39 0.716 1.44
Gaussian ($\sigma=0.75$) GC Scratch 1 epoch 19.58 0.852 $0.15\times 10^{3}$ 21.35 0.875 $0.10\times 10^{3}$ 15.97 0.475 $0.57\times 10^{2}$
50 epochs 23.35 0.893 $7.30\times 10^{3}$ 28.48 0.944 $5.34\times 10^{3}$ 21.67 0.632 $2.88\times 10^{3}$
100 epochs 24.21 0.908 $1.45\times 10^{4}$ 29.47 0.951 $1.08\times 10^{4}$ 23.33 0.651 $5.82\times 10^{3}$
300 epochs 24.65 0.914 $4.09\times 10^{4}$ 29.53 0.954 $3.20\times 10^{4}$ 24.47 0.675 $1.74\times 10^{4}$
PudNet 24.30 0.893 1.49 28.43 0.947 1.17 23.83 0.636 1.45
Finetuning Predicted Parameters. Since a typical strategy for applying a pretrained model to a new dataset is to finetune the model, we evaluate our method by finetuning the models obtained from both the intra-dataset and the inter-dataset settings. To facilitate this evaluation, we incorporate an additional linear classification layer into PudNet, as well as and baselines (excluding ‘Adam Scratch’ and ‘GC Scratch’). Subsequently, we randomly select 10,000 images from CIFAR-100 and 800 images from DTD for finetuning the respective models, while the remaining images are reserved for testing. The results displayed in Table VIII demonstrate that our method outperforms other competitors and achieves the best performance. This indicates that the predicted parameters generated by PudNet can effectively serve as a pretrained model.
Effect of Different Groups of Datasets. We analyze the effect of different groups of datasets on training PudNet. Fig. 5 reports the results obtained by exploiting PudNet to predict parameters for ResNet-18. We construct varying groups of datasets using Fashion-set, CIFAR-100-set, and ImageNet-100-set, respectively. We find that with more groups of datasets for training, our PudNet achieves better performance. This improvement can be attributed to the fact that a larger number of groups provide more opportunities for our PudNet to learn the hyper-mapping relation, thereby enhancing its generalization ability. However, as the number of groups becomes large, the performance increase becomes slow.
@@
Parameter Sensitive Analysis. We analyze the effect of different values of the hyper-parameter $\eta$. Recall that $\eta$ controls the percent of dataset complementary information in the initial residual connection. Fig. 6(a)(b)(c) show the results in terms of ResNet-18. We observe that our model obtain better performance when $\eta>0$ in general. Additionally, our method is not sensitive to $\eta$ in a relatively large range.
@@
Convergence Analysis. We discuss the convergence property of the proposed method by plotting the loss curves with increasing iteration. Here we utilize PudNet to predict parameters for ResNet-18, based on Fashion-set, CIFAR-100-set and ImageNet-100-set respectively. As shown in Fig. 7, the training metric-based loss (abbreviated as metric loss) and training total loss first decrease rapidly as the number of iterations increases, and then gradually decrease to convergence.
TABLE XI: Results of different methods in the Inter-dataset setting for image denoising
[…]
PNSR(dB) SSIM time(sec.) PNSR(dB) SSIM time(sec.) PNSR(dB) SSIM time(sec.)
Gaussian ($\sigma=0.25$) GC Scratch 1 epoch 26.88 0.933 $0.12\times 10^{3}$ 19.89 0.826 $0.71\times 10^{2}$ 13.69 0.443 $0.16\times 10^{2}$
50 epochs 32.24 0.963 $5.82\times 10^{3}$ 25.92 0.913 $3.57\times 10^{3}$ 28.29 0.921 $0.84\times 10^{3}$
100 epochs 34.13 0.973 $1.19\times 10^{4}$ 26.75 0.924 $7.25\times 10^{3}$ 30.55 0.933 $1.69\times 10^{3}$
300 epochs 35.65 0.982 $2.99\times 10^{4}$ 27.61 0.935 $2.21\times 10^{4}$ 32.84 0.947 $5.07\times 10^{3}$
PudNet 33.61 0.971 1.02 26.24 0.920 1.53 30.80 0.931 0.67
Gaussian ($\sigma=0.50$) GC Scratch 1 epoch 26.09 0.912 $0.12\times 10^{3}$ 19.14 0.794 $0.72\times 10^{2}$ 14.08 0.439 $0.17\times 10^{2}$
50 epochs 30.42 0.951 $5.83\times 10^{3}$ 23.81 0.875 $3.58\times 10^{3}$ 25.44 0.831 $0.85\times 10^{3}$
100 epochs 31.26 0.956 $1.19\times 10^{4}$ 24.57 0.888 $7.25\times 10^{3}$ 26.50 0.841 $1.70\times 10^{3}$
300 epochs 32.27 0.964 $3.00\times 10^{4}$ 25.14 0.902 $2.21\times 10^{4}$ 28.59 0.865 $5.07\times 10^{3}$
PudNet 30.48 0.950 1.03 24.10 0.881 1.53 28.20 0.868 0.68
Gaussian ($\sigma=0.75$) GC Scratch 1 epoch 24.41 0.887 $0.12\times 10^{3}$ 18.29 0.763 $0.74\times 10^{2}$ 14.12 0.369 $0.17\times 10^{2}$
50 epochs 28.75 0.934 $5.83\times 10^{3}$ 22.41 0.844 $3.59\times 10^{3}$ 22.73 0.750 $0.85\times 10^{3}$
100 epochs 29.17 0.940 $1.20\times 10^{4}$ 23.09 0.854 $7.26\times 10^{3}$ 23.75 0.771 $1.70\times 10^{3}$
300 epochs 29.92 0.944 $3.00\times 10^{4}$ 23.60 0.868 $2.21\times 10^{4}$ 25.83 0.801 $5.07\times 10^{3}$
PudNet 28.72 0.931 1.03 22.66 0.846 1.53 26.02 0.807 0.68
[…]
Figure 8: Visualizations of testing results on different datasets. The first row displays the noisy images with a Gaussian noise level of $\sigma=0.25$. The second row showcases the denoised images generated by the PudNet using a single forward propagation.
IV-D Image Denoising Experiments
We additionally conduct experiments on a typical image denoising task in the image processing field. Here, we leverage PudNet to generate parameters for a well-established convolutional blind-spot network [65], which is built upon the UNet [66] architecture.
[…]
Performance Analysis of Intra-dataset Image Denoising. Table IX presents the results of image denoising in the intra-dataset setting. To demonstrate the time our method could save, we provide the time of training the blind-spot network from scratch using the widely-used optimizer Adam and the training acceleration technique, GC. We find that it takes around 100, 50, and 50 epochs to train the blind-spot network using the accelerated method GC and the network obtains PSNR values of 28.64 dB, 25.95 dB, and 24.64 dB, as well as SSIM values of 0.953, 0.919, and 0.892 for three different types of Gaussian noise, respectively. In contrast, our method costs less than 1 GPU second to predict the parameters of the blind-spot network while still achieving the comparable performance (PSNR of 28.53 dB, 25.97 dB, and 24.52 dB, and SSIM of 0.950, 0.921, and 0.892) on the three datasets respectively, at least $6,000$ times faster than the training accelerated method.
[…]
V Conclusion and Future Works